1. A rock is dropped from a height of 1.6 meters. How fast is the rock falling just before it hits the ground?


1
Expert's answer
2022-02-11T08:46:21-0500

Given:

"h=1.6\:\rm m"

"g=9.8\:\rm m/s^2"


The law of conservation of energy says

"mgh=\frac{mv^2}{2}"

Hence, the final velocity

"v=\sqrt{2gh}=\sqrt{2*9.8*1.6}=5.6\:\rm m/s"